|Allsky1 System Log|
2018 July 02 - SBIG ST237A camera system has failed and the system has been taken offline
2018June22 - allsky dome cleaned inside and outside and replaced. It was removed;
repaired midnights script and regenerated
2018April04 - The roof of the SCGObservatory blew off in a windstorm during the daytime. Max gusts reached 66kph on our SCG Weather Station. The Allsky1 computer and control system was exposed to a few inches of snow after that, and may or may not be online until the situation is resolved.
2017Dec10 - upgraded the 5vdc power source on the dome heater to 12vdc, resulting in a change of wattage delivered from 0.5 watts to 2.88 watts.
was: v=ir; i=v/r=5/50=0.1 amp; W=IV=0.1*5=0.5 watts
is: i=v/r=12/50=0.24 amps; W=iV=0.24*12=2.88 watts; which is about a x6 increase in heat delivered.
2017Nov29 - The 0.5 watts of heat are not nearly enough to combat snow and ice, although it has been doing not a bad job on condensation. Will ramp that up to a 12vdc supply in the near future. The setup photo at the top of the page has been updated. The black plastic around the dome is a screen to block local lighting.
2017Nov03- computer hardware as of today: Microsoft Windows XP Service Pack 3, on a Dell Precision workstation 370, Pentium 4 CPU @ 2.8GHz with 2GB DDR2 RAM, 250GB SATA drive. The camera controller uses the LPT port and is turned off and on each day via a mechanical timer on its power.
It is connected with a Gigabit network ethernet. The video display resolution is 1024x768. The CCDOps software is run by a mouse match script, so positioning of the CCDOps windows is critical.
2017Oct 22 - added 0.5 watts of heat inside the 4" dome in the way of a 5vdc powered two 100ohm resistors in parallel (50ohm). No noticable difference so far, after one night. Also removed a lot of bad silicon and replaced with sticky velcro
2017March28 - cleaned dome surface with concentrated rainx so repel water better. Dome has adhered to camera but there are still gaps. No heater has yet been installed.
2017Mar20 - system restarted with new small 4" dome overtop of SBIGST237A camera system. Seal not complete, open to air but not rain. Main reason is to protect from rain. Reset start and stop times and noticed frost in the morning. May need heater again as well.
2017Feb27 - water in the lens. The system will be down for the new few days while the lens is removed and dried out.
2017Jan01 - celebrating 10 complete years of this system in operation (since 2006 December). There have been a few downtimes, mainly when the old computers with an LPT: port dies in the outdoor cold (has gotten down to -35C) and the time needed to find a replacement. The camera itself has been solid.
2016Oct20 - looks like the mechanical timer that starts and stops power to the camera control box is failing... the camera shut down at 01:28 EDT this morning instead of 08:00 or so. will remove the timer for now and leave the camera controller on 24/7
2016march31 - fixed code generating daily .avi summary of images.. we including older images as well
2016Jan22 - update code references from 2015 to 2016, especially in the processing of grabbing and displaying midnight images.
2015Oct19 - system failed last night, first time in many months. will manually watch over it this evening. Since the camera was removed from the housing, the CCD temperature and peltier cooler are working much better now, on the order of 25deg C
2015April12 - completely removed camera system from existing housing and 14" dome and mounted in a similar orientation on the same post just the camera system with integrated lens. DOcumentation suggested the system was waterproof and it would remove the huge heating requirements of a 14" dome airspace, a 2nd optical surface and fewer bugs sitting on the inside of the dome.
2015Feb18 - .avis have been small but poor quality.. change starting tomorrow using modified conversion parameters /web/allsky1/ffmpeg -r 10 -i %05d.jpg -vcodec mpeg4 -b 800k today.avi
will increase file size from approx 1.5MB to 4.6MB
2014Dec11 - added new FFMPEG to software resources and from today onward will generate much smaller .AVI files (<3MB) replacing the larger .GIF files (!40MB)
2014Dec10 - removed radiojove software from this computer and put onto standalone laptop
2014Jan25 - added radiojove software radio skypipe to this computer as well as the RJ webcamera. 24vdc power supply for heaters still working. Extreme cold temperature mean the heaters cannot keep up. Exposure time still 90sec. Considering a move to 60 sec. would have greater percentage of dead time as images download, from 10% to 15%. Would get better exposures when moon is up. Moving object detection/exposure should not change.
2013Oct - 12vdc power supply for circulation fans failed.
2013Aug10 - Power system reconfigured, now all parts of system now on UPS.
2013Aug08 - some part of the system is not on the UPS and whenever there is a short power outage, the recording of images for the night ends and does not restart. Flies and bees still getting inside dome. TIme to think about redesigning and rebuilding the camera housing.
2013Feb25 - replaced laptop with desktop again.
2013Feb24 - replaced temp desktop with 10-13 year old dell inspiron laptop (pentium mobile @2.0hz, 512mb RAM, 40gb IDE drive). Totally underpowered and failed to run the ccdsoft software.
2013Jan23 - computer crashed again. turned off around 22:00 today for repair.
2013Jan22 - discovered allsky desktop computer crashed around 23:00 EST Jan 21 due to extreme cold. Restarted at 20:00. Crashed again around 21:08 Jan22.
missed a good meteorite over NY state.
2012Nov21 - the batch script to start the ccdops program under windows 7 now works much better and is automated again.
2012Nov20 - autostart scripts failing the last two weeks.. needed manually startup every single night. Finally think this is now fixed for tonights run.
2012Nov16 - added in American Meteor Society maps for the events we captured.
2012Oct04 - everything seems to be working fine. annotations are good. Built winter insulating cover for desktop allsky1
2012Sep03 - replaced laptop with old desktop for now. mostly operational. needed a lot of configuration tweaking. Also seem to have an older version of CCDSoft then we used to. strange. internal FITS headers are different and that broke some of the scripts. The temperature and exposure data annotations are still failing on the midnight images.
2012Aug31 - allsky1 laptop hard drive fails. Ran diagnostic and many many sectors are toast and unrecoverable. Looking to rebuild a new computer. Allsky1 will be out of service until this is done.
2012Aug24 - updated annotation scripts again to add Eastern Time (-5/-4 UT). Black plastic mask around lens positioned well to block local houses direct light. Starting to shift start and end times again as sunrise and sunset are later and the nights are getting longer. Housing still needs new coat of white paint, dome needs a cleaning.
2012Aug09 - the new png scripts worked well but the uploads to this server did not. Also the noise moves around a lot.. changing to -noise 3 for tonight.
2012Aug08 - change batchfits processes to create .PNG images instead of .JPG; also change -noise 1 to -noise 2 in an attempt to get rid of the noise in the images since the ImageMagick upgrade on June 23, 2012
2012july06 - new network cable laid to observatory and new gigabit switch. changed processing for 'collection of midnights'
2012may01 - network to observatory failure again from april 30th 17:00 to 06:30 today. temporarily fixed, awaiting new network cable installation. Partial data collected. Adding enhanced annotation to images, part 1 (north,south,east,west)
2012april26 - added Aurora Collection
2012april23 - created new collection of images every midnight for an easily accessible historical record of that nights sky.
2012april23 - big wind and snowstorm overnight. Looks like the lens mask has shifted position twice through the storm.
2012march16 - network switch port burned out/dead, moved to a vacant port, all good.
2012march15 - very close lightning storm strikes, lost network to observatory
2012mar07 - data processing glitch.. will take another day to work itself through
2012Feb20 - laptop back in service, allsky partial run, complete run Tuesday Feb 21st evening.
2012Feb17 - laptop harddrive fails, allsky out of service
2012Feb08 - removed internal mask #2, installed slightly larger internal mask #3
2012Feb05 - removed internal mask #1, installed internal mask #2
2012Feb04 - removed external light mask from dome, installed internal mask #1
2012Jan19 - indications of failing network cable to observatory, especially in cold (-20) weather. Will wait for good connection on warm day, transfer over images from local
observatory storage and reprocess manually.
2012Jan11 - removed .mpg generation of large .gif as it was overwhelming the old slow processor on the web server and locking up the whole site for an hour each day.
total combined parallel heater resistance=33ohms@24vdc giving 17 watts of heat.
2011Dec16 - broke the main processing script on Dec 13 and just repaired this morning
2011Dec04 - replaced the dome and platform, replaced circulation fan with a larger one and mounted it inside the dome, remounted the two resistive heaters.
2011Nov04 - fixing some of the code for archiving images and animations.
2011Sep28 - added new "slide show" method of displaying the nights run of images, in addition to keeping the old one.
2011Sep19 - now annotating CCD Temp in Deg C, and now exposure time! CCD temperature starts out at about 17- -11=28degC below ambient. By the end of the nights run it is 12- -11=23degC. This has improved slightly since the housing vent was installed.
The first image of the night shows the camera at ambient.
2011Sep18 - removed housing door and installed vent with insect screen as camera was often 5-10 degrees C above ambient at the start of the nights run.
2011Sep16 - change startccd time from 00:30 UT to 00:01 UT
2011Aug02 - problems on the last two nights runs. should be fixed for tonight... just as the clouds come in :)
2011Jun19 - close to solstice, changed start times to 01:59UT (21:59EDT) as camera is oversaturating and process the last images at 08:00UT (04;00EDT)
2011May13 - startCCD time has been 01:19 UTC (21:19 EDT) and stop time 08:23 UTC (04:23 EDT) for some time now and even then the start and end are overexposing.
Problems with the parallel skypipe software interfering with ccdops seem to have been resolved as well.
2011Mar26 - changed StartCCD time to 00:30 UTC (20:30 EDT) and stop time to 10:00 UTC (06:00 EDT)
2011Feb22 - changed StartCCD time to 23:45 UTC (18:45 EST) and stop time to 11:15 UTC (05:15 EST)
2011Feb15 - changed processing scripts a bit to lighten the images as they had become too dark since migrating over to the laptop.
Changed startccd time to 18:32 EST changed STOP time to 06:04EST
2011Jan13 - another laptop (7 years old) installed with all software and allsky now up and running again.
20110907 - applied coat of rainx to dome. looks pretty scummy and full of dead wasps :) Placed laptop and power supply into environmental box for the winter.